OpenStudy (umulas):

What is the solution to the equation log 2 x – log 2 3 = 2 ?

OpenStudy (anonymous):

is it log 2x or is it log x to the base 2, and also log 23 or log 3 to base 2

OpenStudy (anonymous):

please retype the question

OpenStudy (anonymous):

\[\huge \log_2 x - \log_2 3 = 2\]this i guess

OpenStudy (umulas):

yes the one the np did

OpenStudy (umulas):

I mean as usual logs i didd x - = 2 3 But the answer is not 6.

OpenStudy (anonymous):

\[\huge \log_2 x = 2+ \log_2 3\]\[\huge 2^{\log_2 x} = 2^{2 + \log_2 3}\]\[\huge x = 2^2 \times 2 ^{\log_2 3}\]\[x = 4 \times 3\]

OpenStudy (umulas):

\_/ what how did you get he log and the 2 together?

OpenStudy (anonymous):

o.O which one, how to type it or how to do it?

OpenStudy (umulas):

how to do it.

OpenStudy (anonymous):

\[\huge \log_2 x = 2 + \log_2 3\]then 2^(everything)\[\huge 2^{\log_2 x} = 2^{2+\log_2 3}\]\[\huge 2^{\log_2 x} = x\]and the other side using a^(m+n) = a^m * a^n \[\huge 2^{2+\log_2 3} = 2^2 \times 2 ^{\log_2 3} = 4 \times 3\]